# First-Order Phase Transition of the Schwinger Model with a Quantum Computer

### 20 Dec 2023 v1

@@ Appendix C Tensor networks
Figure 10: Electric field density $\mathcal{F}/g$ against $l_{0}$ for Wilson fermions using no mass shift at $m_{\text{r}}/g=m_{\text{lat}}/g=0.01$. The MSE is calculated as in Fig. 6 and found here to be $7.939\cdot 10^{-4}$.
Appendix D Lattice effects on the phase transition
[…]
$E_{l_{0}<l_{0}^{*}}=a(N-1)g^{2}\frac{l_{0}^{2}}{2},$ (27)
$E_{l_{0}>l_{0}^{*}}=a(N-1)g^{2}\frac{(l_{0}-1)^{2}}{2}+2m_{r}.$ (28)
$l_{0}^{*}=2\frac{m_{\text{r}}}{g}\frac{\sqrt{x}}{N-1}+\frac{1}{2}.$ (29)
In the limit of large $N$, Eq. (29) becomes
$l_{0}^{*}\approx 2\frac{m_{\text{r}}}{g}\frac{\sqrt{x}}{N}+\frac{1}{2},$ (30)
The mass shift decreases with decreasing $ag$ [26, 51]. Further, the factor $1/(1-ag)$ decreases as $ag\to 0$, which leads to an overall decrease of the first term and Eq. (31) tends towards the continuum limit. Hence, we see that the transition point on finite lattices $l_{0}^{*}$ is generally shifted towars values larger than $1/2$ and gets closer to the continuum theory prediction for decreasing lattice spacing, as observed in Sec. IV.2 and supported by Fig. 3.

### 11 Mar 2024 v2

@@ Appendix D Lattice effects on the phase transition
Acknowledgements.
This work is partly funded by: the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie COFUND scheme with grant agreement no. 101034267, the European Union’s Horizon Europe Framework Programme (HORIZON) under the ERA Chair scheme with grant agreement no. 101087126 and the Ministry of Science, Research and Culture of the State of Brandenburg within the Centre for Quantum Technologies and Applications (CQTA). P.N. acknowledges financial support from the Cyprus Research and Innovation Foundation under projects “Future-proofing Scientific Applications for the Supercomputers of Tomorrow (FAST)”, contract no. COMPLEMENTARY/0916/0048, and “Quantum Computing for Lattice Gauge Theories (QC4LGT)”, contract no. EXCELLENCE/0421/0019. A.C. is supported in part by the Helmholtz Association —“Innopool Project Variational Quantum Computer Simulations (VQCS).” The authors also thank the QC4HEP Working Group for discussions. IBM, the IBM logo, and ibm.com are trademarks of International Business Machines Corp., registered in many jurisdictions worldwide. Other product and service names might be trademarks of IBM or other companies.
AUTHOR CONTRIBUTIONS
T.A. drafted the paper, led the calculations for Wilson fermions and performed the MPS computations. P.N. led the calculations for staggered fermions. D.W. advised on hardware experiments with error mitigation. Everyone has contributed to providing directions and ideas and writing the paper.
COMPETING INTERESTS
All authors declare no financial or non-financial competing interests.